OpenStudy (anonymous):

hey all help me plz find the tangent line to f(x)=15-2x^2 at x=1

OpenStudy (radar):

The point would be (1,13) solving for when x=1. Slope would be -4 from taking derivative of f(x). You now have slope and a point. You have the info write the equation for the tangent line. Can you take it from here?

OpenStudy (anonymous):

i attachments you > i want solve Question No. 4 if u can

OpenStudy (radar):

The B Part, as A. was what was above.

OpenStudy (anonymous):

oky the part A. solve it f(x)=15-2x^2 at x=1 f(x)=15-2(1)^2 f(x)=15-2 f(x)= 13 The point would be (1,13) solving for when x=1. Slope would be -4 from taking derivative of f(x). thats true ? and how can Draw The Graph and Part C and D ? can u solve it plz

OpenStudy (radar):

B.\[-(x ^{3}-8)\over x ^{2}-4\] Please note, that the numerator is the difference of two perfect cubes, and the denominator is the difference of two perfect squares. These a special cases and can be factored as follows:\[-(x-2)(x ^{2}+2x+4)\over (x+2)(x-2)\]Now you can do some canceling and get:\[-(x ^{2}+2x+4 \over x+2\]The limit when x approaches 2 would be:\[12\over4\] or 3. B. is 3

OpenStudy (radar):

Excuse me I forgot to consider the minus sign. -3 is the answer

OpenStudy (anonymous):

oky thank you so so much oky the part A. solve it f(x)=15-2x^2 at x=1 f(x)=15-2(1)^2 f(x)=15-2 f(x)= 13 The point would be (1,13) solving for when x=1. Slope would be -4 from taking derivative of f(x). thats true ? and how can Draw The Graph and Part C and D ? can u solve it plz

OpenStudy (radar):

C. I don't understand the problem, just trying to figure it out i. For a = -3 there is a break and f(a) becomes 4 ii. For a =-1, f(a) breaks and becomes a 3 iii. For a= 2, f(a) does not exist iv. For a = 4, f(a)=5 Really am treading in unfamiliar waters here, so these are just SWAGslol Hi robtobey, thigs ok in sunny CA

OpenStudy (anonymous):

ummmm oky thx for u but how can i draw the tangent line for part A (q.4) and do u have solve the part D ( q.4) am sorry

OpenStudy (radar):

The line would be: -4=(13-y)/(1-x) -13y=4x-9 y=(-4/13)x+9/13